The area of a right triangle is 250 in^2. Find the lengths of its legs if one leg is 5 inches longer than the other. (use the five step method.
Please help me

Answer by checkley77(12569)

Area=bh/2
250=x(x+5)/2
250=(x^2+5x)/2
x^2+5x=2*250
x^2+5x=500
x^2+5x-500=0
(x+25)(x-20)=0
x-20=0
x=20 ans. for the base.
20+5 ans. for the height.
Proof:
250=20*25/2
250=500/2
250=250